# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ III Non-equilibrium
[…]
@@
With the lesser Green’s function at hand, lattice observables are obtained as
[…]
We end this section by noting that in order to take the $t_{C}\rightarrow 0$ limit, effectively reinstating equilibrium in the conductor region, the self-energy $\Sigma^{R}$ should be replaced by the usual $i\eta$ factor. Eq. (14) then reduces to the usual $G^{<}(\omega)=iA(\omega)n_{F}(\omega-\mu)$ where $\mu$ is the chemical potential of the conductor region.
@@ IV Results and discussion
With the non-equilibrium formalism in place, we can now circumvent the ambiguities associated with the OAM current by focusing instead on the local intra-atomic OAM density $\langle L_{l}\rangle$ emerging under applied bias. Before moving on to the results, it is worth highlighting why doing so is advantageous. First, the OAM density is an unambiguous observable and directly measurable by using the MOKE [19, 20], unlike the OAM current. Second, the OAM current in not conserved, so while we will discuss $\sigma_{ij}^{l}$ from Eq. (4), we emphasize that the conductivity of a non-conserved current is conceptually problematic. Much of the work on the orbital Hall effect relies on calculating $\sigma_{xy}^{z}$ via the Kubo formula, using it as a figure of merit. However, since the OAM current is not conserved, it is unclear to what extent a nonzero $\sigma^{l}$ will give rise to an orbital accumulation. By instead calculating the OAM density, we are able to directly observe, in real space, the orbital accumulation induced by an electric field. Moreover, we can compare the actual OAM response with that predicted by computing only the orbital conductivity. We note in passing that if an orbital Rashba interaction is added to our model, we find that an orbital diode effect appears which is discussed in more detail in App. C
@@
Figure 4: The expectation value for the OAM operators $L_{i,l}$ at each lattice site $i$ for a system with $N\times N$ sites, $N=40$. The system is governed by Hamiltonian in Eq. (2) with parameters $\mu=-2.85$, $t_{m}=0.1$, $t_{\gamma}=0.5$, $T=0.01$. In all panels, a voltage bias is applied $V=0.01$ and the black arrows signify the direction of the resulting charge current $\vec{j}_{c}$. In (a) and (b), the bias is along one direction: $V_{x}=V$ in (a) and $V_{y}=V$ in (b), while in panel (c) biases are applied in both directions simultaneously: $V_{x}=V_{y}=V$.
@@
where $v_{j}(\boldsymbol{k})=\partial_{k_{j}}H(\boldsymbol{k})$ and $J^{l}_{j}(\boldsymbol{k})=\{v_{j}(\boldsymbol{k}),\tilde{L}_{l}\}/2$. Numerically diagonalizing Eq. (3) allows straightforward calculation of Eqs. (18) and (19). As, our work mainly pertains to real-space imaging of the OAM density, we will omit the quantitative results for $\sigma_{ij}^{l}$ and only state the symmetries it exhibits, relating this to the OAM density. We find that for all values of $t_{\gamma}$ and $t_{m}$, Eq. (18) yields $\sigma_{ij}^{l}=0$ when $i\neq j$, consistent with Fig. 4 (a). Moreover, the orbital conductivity tensor obeys the following symmetry relations
@@
These spatial coordinates symmetries of $\sigma_{ii}$ are unsurprising, as the fourfold symmetry of the Hamiltonian should be reflected in $\sigma_{xx}^{l}$ and $\sigma_{yy}^{l}$. The relative minus signs in Eq. (20), however, are non-trivial. To confirm their existence, in Fig. 4 (b) we consider instead a nonzero bias in the $y$ direction. The resulting OAM response is consistent with Eq. (20); charge current in the positive $y$ direction induces an OAM response with opposite sign relative the current direction compared to Fig. 4 (a).
As we will now demonstrate, the orbital Hall response is only zero when the charge current is applied in a direction aligning with the crystal axes. In Fig. 4 (c), both a vertical and horizontal bias is applied, clearly resulting in a transverse OAM response to the charge current, and is thus an orbital Hall response. Moreover $\langle L_{z}\rangle$ is zero along the direction of the current. Both of these points can be understood by examining how the orbital conductivity transforms under rotation around the $z$ axis. In the original coordinate system, the orbital conductivity tensor takes the form $\sigma^{l}=\text{diag}(\sigma_{xx}^{l},\sigma_{yy}^{l})$. The transformed conductivity tensor $\sigma^{\prime l}$ is then equal to
@@
where $R(\theta)$ is the conventional 2D rotation matrix with angle $\theta$. Since $\sigma_{xx}^{z}=-\sigma_{yy}^{z}$, it follows directly from Eq. (21) that in a coordinate system rotated by 45 degrees relative to the crystal axes, only a Hall response remains for the $z$ component of the OAM. The longitudinal orbital conductivity vanishes, consistent with the behavior we see in Fig. 4 (c). However, because $\sigma_{xx}^{x/y}$ and $-\sigma_{yy}^{x/y}$ are not generally equal, a similar cancellation does not occur for $\langle L_{x}\rangle$ and $\langle L_{y}\rangle$. As a result, they exhibit both longitudinal and transverse responses in Fig. 4 (c), with the transverse response being the dominant one.
The orbital conductivity closely resemble the spin conductivity in altermagnets. With suitable alignment with the crystal axis, their longitudinal components in the $x$ direction are opposite to the components in the $y$ direction. The transverse response to an electric current is therefore the strongest when an electric field is applied in the direction of the band crossings of altermagents, or in our system, the avoided crossings. Thus, the effect in Fig. 4 depends on the orientation relative to the crystal axis, suggesting a comparison to the spin splitter effect in altermagnets [54, 63]. What we observe here is the orbital counterpart, namely the orbital splitter effect. States with positive OAM is deflected in one direction, negative OAM in the other, with the effect being dictated by the crystal orientation.
Despite the similarities in the splitter effects, the microscopic mechanisms giving rise to the spin and orbital polarizations are different. The orbital splitter effect presented here arises with the orbital texture, which only exists out of equilibrium, while the spin polarization in altermagnets follow directly from their compensated spin bands [54] which are polarized even without any current bias. Still, the orbital splitter effect is strongly related to the angle between the charge current and the crystal axis, justifying comparison to the spin splitter effect. Observing it in experiments would require high quality single crystals in order to avoid that the effect is averaged out.
@@
Figure 5: The energy bands for $H(\boldsymbol{k})$ in Eq. (3) with $t_{m}=0.2$ and $\gamma=0.5$ along high-symmetry points. Inset: Showing the small $\boldsymbol{k}$ behavior near $\Gamma$ with dashed lines showing the two values for $\mu$ used in Fig. 6.
@@
Figure 6: The same as in Fig. 4, with only a horizontal voltage $V_{x}=V=0.01$. Otherwise the system parameters are $t_{m}=0.2$, $t_{\gamma}=0.5$, $N=40$, $T=0.01$. In panel (a) $\mu=-3.05$ and in panel (b) $\mu=-2.85$. The black arrows signify the direction of the charge current $\vec{j}_{c}$.
@@
The longitudinal oscillations in Fig. 6 (b) often feature in systems when two or more bands are filled and the system is far below half-filling. Similar oscillations were observed previously in Ref. [31] using the quantum Boltzmann formalism, which demonstrated that local injection of OAM results in an oscillatory orbital torsion and OAM as a function of position. To examine whether these oscillations are tied to the system size, we vary $N_{x}$ in Fig. 7 (b), keeping all other parameters fixed. Summing $\langle L_{y}\rangle$ across rows, we plot the resulting sums as a function of horizontal position. While there are differences between measurement series, all exhibit periodic variation with approximately the same oscillation length, indicating that the oscillation length depends on microscopic parameters rather than system size. To explore this dependence further, we vary $t_{\gamma}$ in Fig. 7 (a). We observe that the oscillation length increases with $t_{\gamma}$. In previous work [36], we found that decreasing $t_{\gamma}$ accelerated the absorption of OAM by the lattice. Here, we observe a similar trend, though in this case, the OAM undergoes sustained oscillations rather than being damped, as in Ref. [36].
@@
The oscillations in the OAM densities in Figs. 6 and 7 is one example demonstrating the lack of OAM conservation. They showcase the limitations of only relying on $\sigma_{ij}^{l}$ to describe OAM transport; Because OAM current is not conserved, $\sigma_{ij}^{l}$ cannot, on its own, accurately describe the OAM behavior in e.g. Fig. 6 (b). This is even more obvious when considering a higher filling fraction than previously, with Fig. 8 showing the induced OAM response. Although some of the effects may be attributed to the Fermi surface being less isotropic at higher filling, it is striking that the OAM response is so intricate, despite the relatively simple system. We stress that our goal is not to provide a microscopic explanation for Fig. 8, but rather to demonstrate that some systems defy simple descriptions of the OAM, making lattice frameworks a viable and helpful alternative to solely calculating the orbital conductivity.
@@
Figure 8: The same as in Fig. 4, with different system parameters. A voltage $V_{x}=V=0.01$ in the $x$ direction is applied, with the black arrow showing the direction of the charge current $\vec{j}_{c}$. Otherwise, the system parameters are $N=40$, $t_{m}=0.1$, $t_{\gamma}=0.3$, $\mu=-2.0$, and $T=0.1$.
We close this section by completing the comparison between the system under consideration and altermagnets. As previously mentioned, $\varepsilon_{2}(\boldsymbol{k})$ and $\varepsilon_{3}(\boldsymbol{k})$ resemble the compensated spin-split bands in altermagnets, and there are strong similarities between the orbital and spin conductivities. However, there are also differences. First, the energy bands of altermagnets maintain the same spin orientation for all $\boldsymbol{k}$, while the orbital characters of $\varepsilon_{2}(\boldsymbol{k})$ and $\varepsilon_{3}(\boldsymbol{k})$ vary significantly with $\boldsymbol{k}$, interchanging near the avoided crossings at $k_{x}=\pm k_{y}$. Unlike in our model, altermagnets can host conserved spin currents and the representation of the spin operators also allow nonzero spin density without any coupling between bands. Moreover, many of the spin transport properties of altermagnets can be inferred straight from their spin band structure [43, 54]. The equivalent OAM properties are concealed, embedded in the orbital texture, making it far more difficult to obtain a full microscopic explanation of the general OAM behavior.
@@ V Conclusions
In this work, we have considered non-equilibrium orbital transport and demonstrated the existence of an orbital splitter effect. OAM accumulation is challenging to predict due to (i) a contribution from multiple bands to the orbital angular momentum accumulation $\langle L_{l}\rangle$, (ii) $\langle L_{l}\rangle$ being nonzero genreally for $l=x,y,z$, and (iii) the nonconservation of the OAM current. Our real-space, tight-binding non-equilibrium framework naturally overcomes the former two, and circumvents the latter by enabling direct imaging of the local OAM density across various systems, rather than relying solely on the orbital conductivity associated with the nonconserved OAM current. The lack of conservation precludes straightforward evaluation of the equation of motions for an OAM operator, as it cannot be written as a continuity equation with the conventional definition of the OAM current. Instead, torque terms emerge [30, 31, 33, 36], allowing the lattice itself to serve as a source or sink of OAM.
Under applied voltage bias, the OAM response in some cases aligns with the predictions of the orbital conductivity. However, in most cases we consider, the OAM density varies considerably in real space due to the torque terms, for example oscillating in the direction of the charge current. Such features cannot be captured by orbital conductivity alone, highlighting the additional insight gained from spatially resolving the OAM. Our approach of doing so, can, in principle, easily be readily extended to other systems, by modifying the tight-binding basis and parameters, along with finding suitable representations of the OAM operators. For instance, it could be applied to the $sp$ system considered in Ref. [10], or used to investigate the OAM properties of specific materials.

### v2

@@ III Non-equilibrium Green function formalism
$=\frac{1}{2\pi i}\int_{-\infty}^{\infty}\mathrm{d}\omega G_{i,j}^{<}(\omega)$ (6)
@@
We use an inverse temperature $\beta=100t^{-1}$ throughout this article, corresponding to a temperature of $T\simeq 116$ K assuming $t$ to be on the order of 1 eV. With the lesser Green’s function at hand, lattice observables are obtained as
$\langle\hat{O}\rangle=\text{Tr}(\rho\hat{O})=\frac{1}{2\pi i}\int_{-\infty}^{\infty}\mathrm{d}\omega\text{ Tr}(\hat{O}G^{<}(\omega))$ (17)
We end this section by noting that in order to take the $t_{C}\rightarrow 0$ limit, effectively reinstating equilibrium in the conductor region, the self-energy $\Sigma^{R}$ should be replaced by the usual $i\eta$ factor. Eq. (14) then reduces to the usual $G^{<}(\omega)=iA(\omega)n_{F}(\omega-\mu)$ where $\mu$ is the chemical potential of the conductor region. We have not considered the equilibrium system in this paper, but have nevertheless included the $i\eta$ factor in addition to the non-zero self-energy contribution from the leads to ensure correct causality. We note, however, that as long as the self-energy contribution from the leads is non-zero, the presence of the $i\eta$ factor does not affect results. $\eta=10^{-10}t$ was used in the calculations.
@@ IV Results and discussion
With the non-equilibrium formalism in place, we can now circumvent the ambiguities associated with the OAM current by focusing instead on the local intra-atomic OAM density $\langle L_{l}\rangle$ emerging under applied bias. Before moving on to the results, it is worth highlighting why doing so is advantageous. First, the OAM density is an unambiguous observable and directly measurable by using the MOKE [19, 20], unlike the OAM current. Second, the OAM current is not conserved, so while we will discuss $\sigma_{ij}^{l}$ from Eq. (4), we emphasize that the conductivity of a non-conserved current is conceptually problematic. Much of the work on the orbital Hall effect relies on calculating $\sigma_{xy}^{z}$ via the Kubo formula, using it as a figure of merit. However, since the OAM current is not conserved, it is unclear to what extent a nonzero $\sigma^{l}$ will give rise to an orbital accumulation. By instead calculating the OAM density, we are able to directly observe, in real space, the orbital accumulation induced by an electric field. Moreover, we can compare the actual OAM response with that predicted by computing only the orbital conductivity. We note in passing that if an orbital Rashba interaction is added to our model, we find that an orbital diode effect appears which is discussed in more detail in App. C.
IV.1 The orbital splitter effect
@@
Figure 5: The expectation value for the OAM operators $L_{i,l}$ at each lattice site $i$ for a system with $N\times N$ sites, $N=40$. The system is governed by the Hamiltonian in Eq. (2) with parameters $\mu=-2.85$, $t_{m}=0.1$, $t_{\gamma}=0.5$, $T=0.01$. In all panels, a voltage bias is applied $V=0.01$ and the black arrows signify the direction of the resulting charge current $\boldsymbol{j}_{c}$. In (a) and (b), the bias is along one direction: $V_{x}=V$ in (a) and $V_{y}=V$ in (b), while in panel (c) biases are applied in both directions simultaneously: $V_{x}=V_{y}=V$.
@@
where $v_{j}(\boldsymbol{k})=\partial_{k_{j}}H(\boldsymbol{k})$ and $J^{l}_{j}(\boldsymbol{k})=\{v_{j}(\boldsymbol{k}),\tilde{L}_{l}\}/2$. Numerically diagonalizing Eq. (3) allows straightforward calculation of Eqs. (18) and (19). As, our work mainly pertains to real-space imaging of the OAM density, we will omit the quantitative results for $\sigma_{ij}^{l}$ and only state the symmetries it exhibits, relating this to the OAM density. In the coordinate system aligned with the crystal axes, we find that for all values of $t_{\gamma}$ and $t_{m}$, Eq. (18) yields $\sigma_{ij}^{l}=0$ when $i\neq j$, consistent with Fig. 5 (a). Moreover, the orbital conductivity tensor obeys the following symmetry relations
@@
The relative minus signs in Eq. (20) are non-trivial. To confirm their existence, in Fig. 5 (b) we consider instead a nonzero bias in the $y$ direction. The resulting OAM response is consistent with Eq. (20); charge current in the positive $y$ direction induces an OAM response with opposite sign relative the current direction compared to Fig. 5 (a).
As we will now demonstrate, the orbital transverse response is only zero when the charge current is applied in a direction aligning with the crystal axes. In Fig. 5 (c), both a vertical and horizontal bias is applied, clearly resulting in a transverse OAM response to the charge current. Moreover $\langle L_{z}\rangle$ is zero along the direction of the current. Both of these points can be understood by examining how the orbital conductivity transforms under rotation around the $z$ axis. In the original coordinate system, the orbital conductivity tensor takes the form $\sigma^{l}=\text{diag}(\sigma_{xx}^{l},\sigma_{yy}^{l})$. The transformed conductivity tensor $\sigma^{\prime l}$ is then equal to
@@
where $R(\theta)$ is the conventional 2D rotation matrix with angle $\theta$. Since $\sigma_{xx}^{z}=-\sigma_{yy}^{z}$, it follows directly from Eq. (21) that in a coordinate system rotated by 45 degrees relative to the crystal axes, only a transverse response remains for the $z$ component of the OAM. The longitudinal orbital conductivity vanishes, consistent with the behavior we see in Fig. 5 (c). However, because $\sigma_{xx}^{x/y}$ and $-\sigma_{yy}^{x/y}$ are not generally equal, a similar cancellation does not occur for $\langle L_{x}\rangle$ and $\langle L_{y}\rangle$. As a result, they exhibit both longitudinal and transverse responses in Fig. 5 (c), with the transverse response being the dominant one.
The orbital conductivity closely resemble the spin conductivity in altermagnets. With suitable alignment with the crystal axis, their longitudinal components in the $x$ direction are opposite to the components in the $y$ direction. The transverse response to an electric current is therefore the strongest when an electric field is applied in the direction of the band crossings of altermagents, or in our system, the avoided crossings. Thus, the effect in Fig. 5 depends on the orientation relative to the crystal axis, suggesting a comparison to the spin splitter effect in altermagnets [62, 71]. The spin-splitter effect has been indirectly observed through torque measurements in RuO2 [72, 73]. What we observe here is the orbital counterpart, namely the orbital splitter effect. States with positive OAM are deflected in one direction, negative OAM in the other, with the effect being dictated by the crystal orientation. Here, the prerequisite for the orbital splitter effect is $t_{m}\neq 0$, requiring breaking mirror symmetries and fourfold symmetry. As mentioned previously, this may be satisfied in oxide heterostructures [47, 48].
Despite the similarities in the splitter effects, the microscopic mechanisms giving rise to the spin and orbital polarizations are different. The orbital splitter effect presented here arises with the orbital texture, which only exists out of equilibrium, while in altermagnets, the spin polarization, as well as their transport properties, follow directly from their compensated spin bands [62], which are polarized even without any current bias. The equivalent OAM properties are concealed, embedded in the orbital texture, making it far more difficult to obtain a full microscopic explanation of the general OAM behavior. Moreover, altermagnets host conserved spin currents, in contrast to the nonconserved OAM currents. Still, the orbital splitter effect is strongly related to the angle between the charge current and the crystal axis, justifying comparison to the spin splitter effect. Observing it in experiments would require high-quality single crystals to avoid the effect being averaged out.
@@ IV.2 OAM oscillations and nonconservation
Figure 6: The same as in Fig. 5, with only a horizontal voltage $V_{x}=V=0.01$. Otherwise the system parameters are $t_{m}=0.2$, $t_{\gamma}=0.5$, $N=40$, $T=0.01$. In panel (a) $\mu=-3.05$ and in panel (b) $\mu=-2.85$. The black arrows signify the direction of the charge current $\boldsymbol{j}_{c}$.
@@
The longitudinal oscillations in Fig. 6 (b) often feature in systems when two or more bands are filled and the system is far below half-filling. Similar oscillations were observed previously in Ref. [36] using the quantum Boltzmann formalism, which demonstrated that local injection of OAM results in an oscillatory orbital torsion and OAM as a function of position. To examine whether these oscillations are tied to the system size, we vary $N_{x}$ in Fig. 7 (b), keeping all other parameters fixed. Summing $\langle L_{y}\rangle$ across rows, we plot the resulting sums as a function of horizontal position. While there are differences between measurement series, all exhibit periodic variation with approximately the same oscillation length, indicating that the oscillation length depends on microscopic parameters rather than system size. To explore this dependence further, we vary $t_{\gamma}$ in Fig. 7 (a). We observe that the oscillation length increases with $t_{\gamma}$, suggesting that the hopping anisotropy determines the period of the oscillations, as it allows for stronger coupling between the OAM and the lattice. This is consistent with previous work [41], where we found that decreasing $t_{\gamma}$ accelerated the absorption of OAM by the lattice. However, in the present case, the OAM undergoes sustained oscillations rather than being damped, as in Ref. [41].
@@
The oscillations in the OAM densities in Figs. 6 and 7 could be related to the lack of OAM current conservation. They showcase the limitations of only relying on $\sigma_{ij}^{l}$ to describe OAM transport; because OAM current is not conserved, $\sigma_{ij}^{l}$ cannot, on its own, accurately describe the OAM behavior in e.g. Fig. 6 (b). This is even more obvious when considering a higher filling fraction than previously, with Fig. 8 showing the induced OAM response. Although some of the effects may be attributed to the Fermi surface being less isotropic at higher filling, it is striking that the OAM response is so intricate, despite the relatively simple system. We stress that our goal is not to provide a microscopic explanation for Fig. 8, but rather to demonstrate that some systems defy simple descriptions of the OAM, making lattice frameworks a viable and helpful alternative to solely calculating the orbital conductivity.
@@
Figure 8: The same as in Fig. 5, with different system parameters. A voltage $V_{x}=V=0.01$ in the $x$ direction is applied, with the black arrow showing the direction of the charge current $\boldsymbol{j}_{c}$. Otherwise, the system parameters are $N=40$, $t_{m}=0.1$, $t_{\gamma}=0.3$, $\mu=-2.0$, and $T=0.1$.
IV.3 Orbital Hall effect
As a last demonstration of our framework, we consider a system with $t_{m}=0$, instead focusing on the orbital Rashba coupling $t_{\text{SO}}$, thus adding $H_{\text{SO}}=t_{\text{SO}}(\sin k_{y}\tilde{L}_{x}-\sin k_{x}\tilde{L}_{y})$ to the Hamiltonian in Eq. (3). $H_{\text{SO}}$ is a natural extension from the original $sp$-model [49] to $d$ orbitals [50]. The OAM responses to a horizontal and diagonal charge current are shown in Fig. 9 (a) and (b), respectively. First, the OHE for $\langle L_{z}\rangle$ can be observed in both panels, showing a genuine transverse OAM response, independent of the current direction, unlike the orbital splitter effect. The OAM is more localized to the system edges compared to the orbital splitter effect, offering another way to disentangle the two effects. We hypothesize that the fairly strong distortions of the OAM in the OHE, also seen in the orbital splitter effect, is related to the nonconservation of OAM current. Second, in both panels, the orbital Edelstein effect can be observed clearly as a bulk value for $\langle\boldsymbol{L}\rangle$ is induced, where the direction of $\boldsymbol{L}$ obeys $\boldsymbol{L}\propto\boldsymbol{z}\times\boldsymbol{j}_{c}$. Figure 9 thus demonstrates that our framework can also reproduce well-known effects, where the spatial resolution may cast them in a new light.
[…]
cc-3.0cmRefer to caption Refer to caption
Figure 9: The same as in Fig. 5, with similar system parameters. The differences are that $t_{m}$ is zero and the orbital Rashba coupling $t_{\text{SO}}=0.1$ is nonzero, otherwise the parameters are the same. The black arrows signify the direction of the charge current $\boldsymbol{j}_{c}$ and varies across the panels.
@@ V Conclusions
In this work, we have considered non-equilibrium orbital transport and demonstrated the existence of an orbital splitter effect. Oxide heterostructures may break the required symmetries for this effect to emerge. OAM accumulation is challenging to predict due to (i) a contribution from multiple bands to the orbital angular momentum accumulation $\langle L_{l}\rangle$, (ii) $\langle L_{l}\rangle$ being nonzero generally for $l=x,y,z$, and (iii) the non-conservation of the OAM current. Our real-space, tight-binding non-equilibrium framework naturally overcomes the former two, and circumvents the latter by enabling direct imaging of the local OAM density across various systems, rather than relying solely on the orbital conductivity associated with the nonconserved OAM current. The lack of conservation precludes straightforward evaluation of the equation of motions for an OAM operator, as it cannot be written as a continuity equation with the conventional definition of the OAM current. Instead, torque terms emerge [35, 36, 38, 41], allowing the lattice itself to serve as a source or sink of OAM.
Under applied voltage bias, the OAM response in some cases aligns with the predictions of the orbital conductivity. However, in most cases we consider, the OAM density varies considerably in real space due to the torque terms, for example oscillating in the direction of the charge current. Such features cannot be captured by orbital conductivity alone, highlighting the additional insight gained from spatially resolving the OAM. Our approach of doing so, can, in principle, be readily extended to other systems, by modifying the tight-binding basis and parameters, along with finding suitable representations of the OAM operators. For instance, it could be applied to the $sp$ system considered in Ref. [10], or used to investigate the OAM properties of specific materials.